# Zhongtian Dong Dep. Physics & Astronomy

### 30 Nov 2023 v1

@@ 5 Conclusion
Our study demonstrates that EQNNs and QNNs outperform their classical counterparts, particularly in scenarios with fewer parameters and smaller training datasets. This highlights the potential of quantum-inspired architectures in resource-constrained settings. The code used for this study is publicly available at https://github.com/ZhongtianD/EQNN/tree/main.
Acknowledgements: KM is supported in part by the U.S. Department of Energy award number DE-SC0022148. KK is supported in parts by US DOE DE-SC0024407. CD is supported in part by College of Liberal Arts and Sciences Research Fund at the University of Kansas.

### 21 Mar 2024 v3

@@ 5 Conclusions
In this paper, we examined the performance of Equivariant Quantum Neural Networks and Quantum Neural Networks, compared against their classical counterparts, Equivariant Neural Networks and Deep Neural Networks, considering two toy examples for a binary classification task. Our study demonstrates that EQNNs and QNNs outperform their classical counterparts, particularly in scenarios with fewer parameters and smaller training datasets. This highlights the potential of quantum-inspired architectures in resource-constrained settings. This point has been emphasized in a similar study recently in Ref. Chang et al. (2023), which showed that an EQNN outperforms the non-equivariant one in terms of generalization power, especially with a small training set size. We note a more significant enhancement in the performance of an EQNN and QNN compared to an ENN and DNN, particularly evident in the anti-symmetric example rather than the symmetric one. This underscores the robustness of quantum algorithms. The code used for this study is publicly available at https://github.com/ZhongtianD/EQNN/tree/main (accessed on 9 March 2024).
While our current study has primarily focused on an EQNN with discrete symmetries, it is crucial to acknowledge the significant role that continuous symmetries, such as Lorentz symmetry or gauge symmetries, play in particle physics. In our future research, we aim to compare an EQNN with continuous symmetries against classical neural networks. Exploring more complex datasets with high-dimensional features is another direction we plan to pursue. However, handling such examples would necessitate an increase in the number of network parameters, prompting an investigation into related issues like overparameterization, barren plateaus, and others.
[…]
Conceptualization, Z.D.; methodology, M.C.C., G.R.D., Z.D., R.T.F., S.G., D.J., K.K., T.M., K.T.M., K.M., and E.B.U.; software, Z.D.; validation, M.C.C., G.R.D., Z.D., R.T.F., T.M., and E.B.U.; formal analysis, Z.D.; investigation, M.C.C., G.R.D., Z.D., R.T.F., T.M., and E.B.U.; resources, Z.D., K.T.M., and K.M.; data curation, G.R.D., S.G., and T.M.; writing—original draft preparation, Z.D.; writing—review and editing, S.G., D.J., K.K., K.T.M., and K.M.; visualization, Z.D.; supervision, S.G., D.J., K.K., K.T.M., and K.M.; project administration, S.G., D.J., K.K., K.T.M., and K.M.; funding acquisition, S.G. All authors have read and agreed to the published version of the manuscript.
[…]
This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0025759. SG is supported in part by the U.S. Department of Energy (DOE) under Award No. DE-SC0012447. KM is supported in part by the U.S. Department of Energy award number DE-SC0022148. KK is supported in part by the US DOE DE-SC0024407. CD is supported in part by the College of Liberal Arts and Sciences Research Fund at the University of Kansas. CD, RF, EU, MCC, and TM were participants in the 2023 Google Summer of Code.
[…]
Not applicable.
[…]
The dataset used in this analysis was sampled from Equations (1), (4) and (6).
[…]
The authors declare no conflicts of interest. The funders had no role in the design of the study; in the collection, analyses, or interpretation of data; in the writing of the manuscript; or in the decision to publish the results.
[…]
Abbreviations The following abbreviations are used in this manuscript:
API Application Processing Interface
AUC Area Under the Curve
DNN Deep Neural Network
ENN Equivariant Neural Network
EQNN Equivariant Quantum Neural Network
HEP High-Energy Physics
LHC Large Hadron Collider
MDPI Multidisciplinary Digital Publishing Institute
ML Machine Learning
NN Neural Network
QML Quantum Machine Learning
QNN Quantum Neural Network
ROC Receiver Operating Characteristic
[…]